Hey I was having trouble with this word problem please help.( The width of a rectangle is 6 kilometers less than twice it's length. If the area is 108 square kilometers, find the dimensions of the rectangle. )How would I set this problem up?
Answer by scott8148(6628)

let length = x

"The width of a rectangle is 6 kilometers less than twice it's length" ___ w = 2x - 6

area = length * width = x * (2x - 6) = 2x^2 - 6x

"If the area is 108 square kilometers" ___ 108 = 2x^2 - 6x ___ 0 = x^2 - 3x - 54

factoring ___ 0 = (x - 9)(x + 6)

so, x = 9

looks like the question switched length an width